The College of Agriculture obtained the following data representing the one-week growth in centimeters of 20 newly planted soybean plants:

1.3 4.9 3.9 0.8 4.1 1.1 3.1 2.2 2.4 2.4

2.4 3.9 1.8 3.9 3.9 4.1 3.9 2.4 4.0 4.2

Find the following:

a. Mean 

b. Median 

c. Mode 

d. Variance 

e. Standard Deviation 

f. Range  


1
Expert's answer
2021-11-16T11:27:57-0500

(a) Mean

"Mean = \frac{\Sigma X}{N}"

"= \frac{(1.3+4.9+3.9+0.8+4.1+1.1+3.1+2.2+2.4+2.4+2.4+3.9+1.8+3.9+3.9+4.1+3.9+2.4+4.0+4.2)}{20}"

"= \frac{60.7}{20}=3.035"


(b) Median

"Median = \frac{(\frac{n}{2})term+(\frac{n}{2}+1)term}{2}"

Firstly need to arrange data in ascending order like below:


"Median = \frac{(\frac{20}{2})term+(\frac{20}{2}+1)term}{2}=\frac{3.1+3.9}{2}"

Median = 3.5

(c) Mode

Mode represents the number of observations in a data set that is occurring most of the time. As shown in above table, 3.9 is occurring 5 times in the data set. This mean mode is 3.9 in case of given data set.


(d) Variance

"s^{2}=\frac{\varSigma (X-Mean)^{2}}{n-1}"

"s^{2}=\frac{27.37}{20-1}=1.44029"


(e) Standard Deviation

"s= \sqrt{Variance}=\sqrt{1.44029}=1.20012"

(f) Range

Range = Maximum - Minimum

Range = 4.2 - 1.3 = 2.9